Status Report

NASA Space Station On-Orbit Status 17 May 2004

By SpaceRef Editor
May 17, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 17 May 2004

All ISS systems continue to function nominally except those noted previously or below.  Underway: Week 3 of Expedition 9 (and a pretty tough one for the new crew).

After station inspection and morning hygiene, before breakfast and first exercise, the ISS crew performed their second session of the periodic Russian biomedical assessments PZEh-MO-8 (body mass measurement) and PZEh-MO-7 (calf volume measurement).  Padalka set up the MO-8 “scales” equipment and later broke it down and stowed it away.

Also before “savtrak” (breakfast) and physical exercise, CDR Padalka and FE/SO Fincke completed their first session of the Russian crew health-monitoring program’s medical assessment MO-9/Biochemical Urinalysis.  CDR stowed the hardware afterwards. [MO-9 is conducted regularly every 30 days (and also before and after EVAs) and is one of five nominal Russian medical tests adopted by NASA for US crewmembers for IMG PHS (Integrated Medical Group/Periodic Health Status) evaluation as part of the “PHS/Without Blood Labs” exam.  The analysis uses the sophisticated in-vitro diagnostic apparatus “Urolux” developed originally for the Mir program.  Afterwards, the data were entered in the medical equipment computer (MEC)’s special IFEP software (In-Flight Examination Program).]

Later in the day, the crew completed their first PHS/Without Blood Labs exam and performed the clinical evaluation, each one acting first as CMO (crew medical officer) and then as the examined subject. Afterwards, Mike Fincke completed data entry/download for both of them, and stowed the PHS hardware.  [The PHS exam is guided by special software (IFEP, in-flight examination program) on the medical equipment computer (MEC).]

Fincke performed another checkup on the EMU batteries charged recently. [Installing first batteries #2029 & #2030 in BCs (battery chargers) 3 and 4, later batteries #2045 & #2046, he conducted OCV (open circuit voltage) tests on each of them to verify their health.  BCs 1 & 2 remain inoperative.]

OCV tests of two PGT (pistol grip tool) batteries were successfully conducted yesterday off the “job jar” task list.  The batteries were installed and charged early today in a new EHIP (EMU helmet interchangeable portable) charger brought up some time ago on the Shuttle.  [Charging of each of the NiMH (nickel metal hydride) batteries took ~180 minutes.]

Later, Mike installed freshly regenerated Metox (metal oxide) CO2 absorption canisters in the two EMU spacesuits intended for the don/doff dry run on Wednesday (5/19). [Canister #20 was installed in suit 3011 (whose liquid cooling system the crew had successfully “degassed” on 5/12), can #21 in EMU 3005.]

The crew also finished final preparations of the Airlock (A/L)’s equipment lock (E/L).  With A/L configuration (complete with unstowed EMU equipment bag and EMU servicing kit) and battery preparations accomplished, both onboard and ground systems are reported ready for the don/doff dry run on Wednesday. [The crew was provided with a detailed outline of the dry run, including a list of “gotchas” (i.e., things to be considered/looked at before to the dry run).  For further clarification, a 20-min. tagup with the ground by S/G (space-to-ground) audio is scheduled for tomorrow.  Also, a 2-hr. teleconference for reviewing all EMU procedures is planned for 6/4, a week before the spacewalk.  Major activities to be exercised during the dry run are (1) A/E E/L preparation; (2) EMU donning/doffing without IV (intravehicular) crewmember present; (3) EMU prebreathe without IV present; and (4) post-EVA activities without IV present.  The session includes mounting and dismounting of SAFER (simplified air for EVA rescue) units, and test of EMU solo doffing, for the cased of one EV crewmember being incapacitated after a spacewalk.  The EMUs will not be purged during the dry run in order to conserve ISS oxygen.]

The word from Russia is that the new U.S. OBS (operational biomedical instrumentation system) has been manifested on the next Progress flight, 14P. [OBS allows Flight Surgeons on the ground to observe the crewmember’s heart rate and ECG (electrocardiogram) trace during the spacewalk.  The OBS hardware presently on board cannot be tested to verify if its current-limiting capability is still active (to prevent electrical shocks).]

Gennady Padalka performed routine maintenance on the Service Module (SM) toilet facilities (ASU), replacing the pretreat container (E-K) plus hose with a new assembly and discarding the old one.  The CDR also completed the regular daily maintenance/inspection of the SM’s SOZh environment control & life support systems. [E-K contains five liters of pre-treat solution, i.e., a mix of H2SO4 (sulfuric acid), CrO3 (chromium oxide, for oxidation and purple color), and H2O (water).  The pre-treat liquid is mixed with water in a dispenser (DKiV) and used for toilet flushing.]

Padalka later worked on the SM electrical power system, removing the failed #6 storage battery and replacing it with the spare 800A.  The ZRU charge/discharge unit #6 was deactivated by TsUP/MCC-Moscow beforehand and later reactivated.  [The old battery was pre-packed for disposal in Progress.  This restores the full set of eight SM batteries to operation.]

The Science Officer unstowed, set up and configured the hardware for the ADUM (Advanced Diagnostic Ultrasound in Microgravity) for the first session scheduled for tomorrow. [Round 1, involving bone scan activities, will be followed by a second session on Thursday, for lower abdominal and retroperitoneal scans.  Preparations today also included configuring the scanheads and clearing space in front of the HRF (Human Research Facility) rack for setting up and fastening the CMRS (Crew Medical Restraint System), not required for tomorrow’s ultrasound scanning activities, but for Thursday’s.]

The CDR meanwhile broke out and set up the equipment for the periodic Russian MO-10 “Hematokrit” testing scheduled for tomorrow, his second.  [MO-10 measures the hematocrit (red cell mass) value of the blood (as a well-known phenomenon of space flight, red blood cell mass {normal range: 30-45%} tends to go down over time).  Tomorrow’s medical tests for the new station CDR also include the MBI-1 SPRUT-K experiment, part of Russian medical research on the distribution and behavior of human body fluids in zero gravity.]

Fincke conducted the regular routine status checkup of autonomously operating Increment 9 payloads.

Mike also performed the daily leak check of the Lab window’s inter-pane space (“Volume D”), using the “Aeolus” scopemeter with pressure probe.  Today’s pressure reading was 518 Torr/mmHg, or 10 psi (still following the trend of a steady leak rate of ~27 Torr/0.52 psi per day from the cabin into Volume D.)

During the crew day, the MCA (major constituents analyzer) was ground-commanded from LEM (life extending mode) to Operate mode for a “zero” calibration and data take, after which it was configured back to LEM to minimize its vacuum chamber’s ion pump operations.

At 12:05pm EDT, the crew participated in a 20-min. educational PAO event with the Maine Space Grant Consortium, Great Salt Bay School, Damariscotta, ME.  Students and teachers from Great Salt Bay School, Damariscotta, and Bristol Consolidated School, Pemaquid, talked with the ISS crewmembers who had prepared themselves with student questions uplinked beforehand.

The crew completed their daily 2.5-hr. physical exercise program on TVIS, RED exerciser and VELO cycle with load trainer.  After reviewing the exercise video recorded on 5/14 on TVIS in motorized mode (which did not reveal anything new of relevance to the ongoing troubleshooting efforts), ground engineers gave the Go for the crew to return to motorized ops when using the treadmill.

The next periodic recharging of the Motorola-9505 Iridium satellite phone is scheduled for tomorrow (5/18).  A revised safety certification has been signed.

Update on EarthKAM participation:  Total number of requests for EarthKAM photos during last week’s session reached 1139 (second highest to date), from 61 schools.

Today’s CEO targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Java (Dynamic event.  Weather may well persist cloud free, with opportunities for panoramas look left and right along the volcanic chain.  Looking for Jakarta just left on the north coast of Java), Khartoum, Sudan (nadir pass over this city.  The entire urban region includes the tent city of Omdurman on the opposite [west] side of the Nile River), Dust events, Central Andes (Dynamic event.  Winter sees an increase in the strength of westerly winds at 14,000-20,000 foot altitudes of the arid Andes [in the fifteen degrees south of Lake Titicaca].  Dust plumes can appear at any time; at the source they cast characteristic shadows.  Longer plumes appear light brown against the downwind cloud cover over Argentina), and High Central Andean Glaciers (looking slightly right for a mapping pass of detailed images of snow and ice-capped volcanoes).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status  (as of today, 2:07pm EDT).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On.  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring.  MCA (major constituents analyzer) is in Life Extending Mode (LEM).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On, SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation failed 4/20).  SFOG slot #2 fan suspect (not usable).
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 25.1; ppO2 (mmHg) — 156.5; ppCO2 (mmHg) — 3.3;
  • SM Transfer Compartment:  Pressure (mmHg) — 751; temperature (deg C) — 20.7.
  • FGB Cabin:  Pressure (mmHg) — 755; temperature (deg C) — 21.1.
  • Node:  Pressure (mmHg) — 748.98; temperature (deg C) — 22.7 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 750.98; temperature (deg C) — 24.6; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 751.18; temperature (deg C) — 25.3; shell heater temp (deg C) — n/a, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • (n/a = data not available)
  • PMA-1:  Shell heater temp (deg C) — n/a
  • PMA-2:  Shell heater temp (deg C) — n/a.

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Directed position (blind, dual-angle mode with “night glider”/”sun slicer” drag reduction sweep, non solar-tracking” 2B @ 235 deg, 4B @ 125 deg.)
  • SM batteries:  The new battery #6 is in Cycle (RTs) mode; all other batteries (7) are on line in “Partial Charge” mode.
  • FGB batteries:  Battery #5 is off line; all other batteries (5) are on line in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
  • GNC-2 MDM is prime; GNC-1 is backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off (backup).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22/03).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22/03).
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

Propulsion System:

  • Total propellant load available: 3689 kg (8133 lb) as of 5/13/04;  [SM(755) + FGB(2934) + Progress M-1(0)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 2 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2’s RPC-17 failed 4/21/04).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH YVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, y-axis in velocity vector [yaw: -90 deg, pitch: -9 deg, roll: 1.7 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally (may require a mask).
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this afternoon, 6:56am EDT [= epoch]):

  • Mean altitude — 360.4 km
  • Apogee — 367.7 km
  • Perigee — 353.1 km
  • Period — 91.75 min.
  • Inclination (to Equator) —  51.6259 deg
  • Eccentricity — 0.0010833
  • Orbits per 24-hr. day — 15.69
  • Mean altitude loss last 24 hours — 70 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 31349
  • For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.